find for X
15in/3/8in=6ft/x

Answer by checkley75(3666)

15/(3/8)=6/x cross multiply
15x=3/8*6
15x=18/8 divide both sides by 12:
x=(18/8)/15
x=18/8*1/15
x=18/120
x=.15 feet is the answer.
proof
15/(3/8)=6/.15
15*.15=6*(3/8)
2.25=18/8
2.25=2.25